Suppose that PX=.25, PZ=.4, P[X|Z] = .3. What is a) P[EUF] b) P[F|E]
Answer #1

Solution:-

P(E) = 0.25

P(F) = 0.40

P( ) = 0.30

a) P(E or F) = 0.53

P(下) =-. P(F)

P(E n F) = 0.40 * 0.30

P(Ecap F) = 0.12

P(E or F) = P(E) + P(F) - P(E and F)

P(E or F) = 0.25 + 0.40 - 0.12

P(E or F) = 0.53

b) P(-) 0.48

P(rac{F}{E})=rac{P(Ecap F)}{P(E)}

P(rac{F}{E})=rac{0.12}{0.25}

P(-) 0.48